OpenStudy (anonymous):

solve for m . √m-7 = n + 3

OpenStudy (lgbasallote):

\[\sqrt m - 7 = n+ 3?\]

OpenStudy (lgbasallote):

ok i'll assume that's what you meant... \[\sqrt m - 7 = n+3\] \[\implies \sqrt m = n + 3 +7\] \[\implies \sqrt m = n + 10\] \[\implies m = (n+10)^2\] simplify this

OpenStudy (lgbasallote):

how ever if you meant \[\sqrt{m-7} = n+3\] \[\implies m - 7 = (n+3)^2\] \[m = (n+3)^2 + 7\] simplify this

OpenStudy (anonymous):

\[\sqrt[3]{-125n^12}\]